Shift Summary for Aug 19 OWL, 1999

Richard Kessler (rkessler@fnal.gov)
Thu, 19 Aug 1999 07:19:06 -0500 (CDT)

Shift Summary for OWL, Thursday Aug 19, 1999

On Shift: R.Kessler and Al Erwin

Hours of beam delivered at all intensities: 2.4
Hours of beam delivered at nominal intensity: 2.4
Hours of E832 data (special runs not included): 2.2
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: 10 minutes
had to ramp NM4AN magnet down/up when beam arrived
because it was tripped by accident during NM2EU repair.

========================================================================
Runs ended this shift:
Run 14265 LASER SCAN
Runs 14266-71 COSMICS
Run 14272 BEAM_832 in progress
========================================================================
Accesses during shift:
one: Al found loose DC1Y-OR cable and plugged it back in.
There was no beam so we didn't lose any beam.

========================================================================
Overview of shift:

NM2EU magnet tripped at start of shift.
After Al got bored he went downstairs to fix DC1Y-4A paddle
that was not plugged in properly.
DC1Y ineff is much less in next BEAM_832 run.

We finally got beam around 5am, but the beam x-position wandered
by +-5 wires for a long time.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14265 00:10 Begin LOOOOOONG LASER_SCAN
00:43 End scan. EXCELLENT!

14266 00:55 Begin COSMICS.

1:56 Web-page photograph of charred NM2EU cables suggest
that the problem has been isolated.

2:30 Run hung up due to DDD1 stuck veto.
2:45 Access for DC1Y cable check. Plugged in loose cable.

14267 3:00 drt_fresh_start and then INIT COSMICS run,
but gateway hangs on ktev1.
Reset DDDs and VME crate, drt_fresh_start.
DAMP process core dumps!
Log-out of da_ktev3, login, then drt_fresh_start again.

14269 3:15 COSMICS run finally started ! And no ktevana errors
at the start of run! Al was reading his DC-OR
scalers throughout the init process ... could that
somehow fix crate 53 ????????
CsI 226 has PMT gain drift of -9%; shift crew has
too much sympathy for Julie and decides not to call
[besides, no official complaints from Val].

14271 3:35 Started another COSMICS run with Al reading his
DC-OR scalers ...
but got lots of ktevana errors this time. Oh well.
CsI 226 gaindrift = -8.6%

3:45 MCR calls saying beam coming soon. Stop run 14271.

INIT BEAM_832 but notice that PTKICK= 1.6 MeV/c ...
MCR says maybe they accidentally tripped our magnet.
They cannot turn on other magnets so beam delayed.

Execute AN832DOWN and then AN832UP to be sure
that NM4AN is set correctly.

4:35 Beam comes with SEM=2.8 E12 ...
begin AN832UP ... MCR didn't warn us about beam.

14272 4:50 BEAM_832 Starts; SEM ~ 5.2E12 and X-position off
target by five wires.
CsI 226 gain drift improves to -6%.

18 Beam position OK at spill 18.

20-22 Beam stop closed for unknown reasons ??

Beam X-position drifting +-5 wires since autotune
doesn't work and MCR is trying manually to tune beam.

7:00 Page Rick Ford to work on stable beam tune.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

- beam X-position stability
- CsI 226